Parth (parthkohli):

If a nine-digit number \(x\) is formed using 1, 2, 3, 4, 5, 6, 7, 8, 9 without repetition, then find the probability that the non-negative difference among all the digits equidistant from both ends is 1.

Parth (parthkohli):

The central number has to be odd. 5 ways to choose an odd number. Now if 3 is chosen then the fixed pairs are (1, 2), (4, 5), (6, 7), (8, 9) so 4! ways to move the pairs around and \(2^4\) ways to flip within the pair.\[\frac{4!\times 5 \times 2^4}{9!}= \frac{1}{189}\]Not a part of the choices...

OpenStudy (thomas5267):

How does this work? 864,152,579 is one such number. But 5 is equidistant to both ends and 5-5=0.

Parth (parthkohli):

Nah, the number at the center is not considered, because that way, there would be no such number.

OpenStudy (firekat97):

what are your options?

ganeshie8 (ganeshie8):

Just to confirm that your answer is correct, I am getting the same 1920 numbers that satisfy the given requirement by bruteforce : https://jsfiddle.net/ganeshie8/Lqbagygh/embedded/result/

Parth (parthkohli):

Oh, that question has an answer that exactly matches mine. Good.

ganeshie8 (ganeshie8):

I don't see 1/189 in that question...

Parth (parthkohli):

Exactly, it doesn't match the choices.

ganeshie8 (ganeshie8):

I'm not gonna spend more time on this.. typoes/gross mistakes are typical in textbooks written by indian authors in rush

Parth (parthkohli):

Ahahaha, why that username

ganeshie8 (ganeshie8):

"non- negative difference" are we interpreting this particular phrase correctly ?

Parth (parthkohli):

I guess we are.

Parth (parthkohli):

\[|a-b| = 1\]

OpenStudy (thomas5267):

Non-negative difference does not make much sense since the order of the subtraction is not specified.